OpenStudy (jiteshmeghwal9):

Harry can complete \(\dfrac{1}{4}\) of a work in 5 days & Joe can complete \(\dfrac{1}{5}\) of the same work in 6 dyas. How long would both of them take to complete the work, if they work together ?

OpenStudy (anonymous):

try this way of thinking: H can run 1/4 per 5 days so \[\frac{ 1 }{ 4 } \div5\]= how long he can run per day

OpenStudy (anonymous):

do you think this inspire you sth? if not we'll keep discuss step by step

OpenStudy (jiteshmeghwal9):

sorry but I'm nt getting it :(

OpenStudy (jiteshmeghwal9):

@ranxu6j3

OpenStudy (anonymous):

that's ok so Harry's speed would be 1/20 per day and Joe 1/30 per day (1/5 div 6)

OpenStudy (anonymous):

can you agree with this?

OpenStudy (jiteshmeghwal9):

yup !

OpenStudy (anonymous):

great in this way, you can assume the whole work =1 then when they work together the speed becomes 1/20 +1/30 =1/12 per day so they can finish 1/12 of the work per day then can you think of it?

OpenStudy (jiteshmeghwal9):

But in how many days will they complete their work together ?

OpenStudy (anonymous):

they can finish 1/12 of the work per day so, 1 div 1/12= 12 days

OpenStudy (jiteshmeghwal9):

\[1 \div \dfrac{1}{12}=12\] ohh! i gt it thanx :)

OpenStudy (anonymous):

OH THAT'S GREAT

OpenStudy (jiteshmeghwal9):

Just a little slip from mind :P

OpenStudy (anonymous):

can you agree with this way of thinking? it doesn't matter, people always slip from mind

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (anonymous):

good ^^

OpenStudy (jiteshmeghwal9):

:)